OpenStudy (anonymous):

[7.02] Choose the correct simplification of (4x3 + 3x2 − 6x) − (10x3 + 3x2).

OpenStudy (ivettef365):

just take out parenthesis and you have 4x^3 +3x^2 -6x -10x^3 -3x^2 now just combine like terms

OpenStudy (anonymous):

which we have to combine x right

OpenStudy (ivettef365):

combine just like terms like 4x^3 -10x^3 = (4-10= -6) therefore that would be -6x^3, now combine the other like term

OpenStudy (anonymous):

3x2+3x2=6x2

OpenStudy (ivettef365):

is actually 3x2 -3x2 = 0 that will cancel

OpenStudy (anonymous):

I for got about the negative in front of the x

OpenStudy (ivettef365):

is ok, so did you get the answer ?

OpenStudy (anonymous):

yep thanks

OpenStudy (ivettef365):

yw :)